2. The hardness of water (hardness count) is usually expressed as parts per million (by mass) of CaCO3. What is the molar concentration of Ca²+ ions in a water sample with a hardness count of 175 ppm?



1
Expert's answer
2021-11-29T23:02:01-0500

1.75 × 10⁻³ M


Explanation:


Given:


Hardness = 175 mg CaCO₃/L


or mass of CaCO₃ in 1 L = 175 × 10⁻³ grams


Now,


Molar mass of CaCO₃ = 40 + 12 + ( 3 × 12 ) = 100 grams/mol


The concentration of CaCO₃ = number of moles / volume


or


= (175 × 10⁻³ / 100) / 1


or


= 1.75 × 10⁻³ M


Now,


CaCO₃ ⇒ Ca²⁺ + CO₃²⁻


1.75 × 10⁻³ M 1.75 × 10⁻³ M 1.75 × 10⁻³ M


thus,


the concentration of Ca²⁺ ions in water = 1.75 × 10⁻³ M